OpenStudy (erinweeks):

Use Cramer's rule to solve the system. 2x + 4y - z = 32 x - 2y + 2z = -5 5x + y + z = 20 A. {( 1, -9, -6)} B. {( 2, 7, 6)} C. {( 9, 6, 9)} D. {( 1, 9, 6)}

OpenStudy (erinweeks):

@ganeshie8

OpenStudy (mandre):

It's multiple choice so does it matter how you solve it? To me Cramer's rule would be too slow to solve a basic problem like that.

OpenStudy (anonymous):

Still good to know how to do it. For Ax=b, Cramer's rule states that \[x _{i}=\det(A _{i)}/\det(A)\] where Ai is the matrix formed by replacing the ith column with b.

OpenStudy (anonymous):

Not seeing a way to show a 3x3 matrix with the equation tools. Let me think how to show this.. lol

OpenStudy (mandre):

It's the bottom selection in the Equation tool. Bottom left option.

OpenStudy (anonymous):

That's a 2x2 unless there's an option to change that somewhere.

OpenStudy (mandre):

\[\left[\begin{matrix}1 & 2 & 5 \\ 3 & 4 & 6 \\ 7 & 8 & 9 \end{matrix}\right]\] 1 & 2 & 5 \\ 3 & 4 & 6 \\ 7 & 8 & 9

OpenStudy (mandre):

\\ for extra row and & for extra columns

OpenStudy (anonymous):

Ah thanks so much.

OpenStudy (mandre):

Pleasure. Go for it lol. I did Cramer's rule last semester but I'm a bit rusty. Will have to review it again.

OpenStudy (anonymous):

Take the determinant of:\[\left[\begin{matrix}32 & 4 & -1\\ -5 & -2 & 2\\20 & 1 & 1 \end{matrix}\right]\]to find the solution for x. Insert the b column matrix into the 2nd column for y and the 3rd for z. As Mandre said though, typically you don't want to do this for 3x3 matrices as it is time consuming. Especially when you have high entries like 32 and 20 lol.

OpenStudy (erinweeks):

i dont get it at all. im sorry

OpenStudy (anonymous):

Do you know how to take the determinant of a 3x3 matrix?

OpenStudy (erinweeks):

nope

OpenStudy (anonymous):

Have you ever computed curl of two 3 dimensional vectors?

OpenStudy (anonymous):

Not curl. Cross product. Sorry wrong term.

OpenStudy (erinweeks):

ive done cross product if your talking like multiplication

OpenStudy (anonymous):

As in the vector operation. Sometimes called outer product. Probably not but that's similar. Um ok let me find an explanation and link it.

OpenStudy (erinweeks):

okay

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Determinant scroll down to 3x3 matrices.

OpenStudy (erinweeks):

alright makes sense thats alot lol

OpenStudy (anonymous):

Yeah it's kind of a lot of work once you get above 2x2 or 3x3 which is why you'd probably use software lol. But the principle is nice to know.

OpenStudy (erinweeks):

okay well can you show me how to solve this so i know for sure

OpenStudy (anonymous):

Okay I'm using Wolfram Alpha to solve the determinants so I can save some time. I'm going to solve all four of the ones we'll use and then you can do that on your own for practice if you want. XD

OpenStudy (erinweeks):

thats fine..

OpenStudy (anonymous):

\[\left[\begin{matrix}32 & 4 & -1 \\ -5 & -2 & 2\\ 20 & 1 & 1\end{matrix}\right]\] Okay the determinant of this is 17. I may walk you through of the determinants just so you can see it but it's just a lot of calculations to do 4 of them.

OpenStudy (anonymous):

And that one is A1 btw. I should label these.

OpenStudy (erinweeks):

dont even worry about it maybe another day

OpenStudy (anonymous):

Are you sure?

OpenStudy (erinweeks):

yea

OpenStudy (anonymous):

Ok good luck with it.

OpenStudy (anonymous):

The answer is D anyhow.